OpenStudy (anonymous):

Solve the fallowing equations 4e-8=32 please show steps??

OpenStudy (anonymous):

r u looking for the answer to e?

OpenStudy (anonymous):

this this the quistion idk??? it says 4e-8=32

OpenStudy (anonymous):

umm well if your looking for e then: 4e - 8 + 8 = 32 + 8 4e = 40 4e/4 = 40/4 e = 10 I dont know what else you could be looking for though :/

OpenStudy (anonymous):

it says solve the salloing equations write the variables above their valures to find the name of a famous volcano? 4e-8=32?

OpenStudy (anonymous):

i have no idea what else it could be >.< sorry rebecca....